# STD-LLM: Understanding Both Spatial and Temporal Properties of Spatial-Temporal Data with LLMs

### v1

@@ 3.2 Model Structure
The node embedding obtained through the eigendecomposition of the Laplacian matrix not only assists the LLM in understanding the topology structure in the spatial-temporal data. Furthermore, since the trainable parameters are independent of the graph, we can easily transfer our model between different graph structures, which is key to our model’s ability to perform well in zero-shot learning.
Figure 1: Model architecture. Condition missing refers to the additional generation of missing data on the training set for the purpose of training. Module A provides spatial-temporal embeddings for node and temporal information. The spatial-temporal tokenizers B and C construct temporal and spatial tokens from different perspectives. Module D builds higher-order spatial tokens based on the spatial tokens, which is the output of C. Module E models the impact of external systems.
[…]
Figure 2: Spatial and Temporal tokens.
[…]
Figure 3: Hypergraph Learning.
Spatial Tokenizer.
The spatial tokenizer is designed to convert the spatial-temporal data into sequential tokens by aggregating the information along the spatial dimension. Specifically, each node’s feature vectors over $P$ time slices are encoded into a spatial token, as shown in Figure 3. We further include the topology structure of nodes into the spatial token by incorporating our proposed topology-aware node embedding. To keep the temporal information in the spatial token, We equip the spatial token with the time embedding, including $\mathbf{E}_{t}\in\mathbb{R}^{d_{t}}$ for the time-of-day, $\mathbf{E}_{w}\in\mathbb{R}^{d_{t}}$ for the day-of-week. It is worth noting that missing data are typically padding with the same value, which can result in high weights between missing positions during attention operation. This may lead to an inability to aggregate valid information, thereby affecting the accuracy of forecasting and imputation. Therefore, it is necessary to enhance the model’s ability to perceive missing values. We generate mask tokens based on the mask matrix to enhance the model’s ability to perceive missing values. The entire computational process in spatial tokenizer can be formalized as follows:
[…]
$\bm{Z}_{S}=\text{LayerNorm}(\mathbf{Z}+\bm{Z}_{mask}),$
where $\mathbf{W}_{*}$ and $\mathbf{b}_{*}$ denote the trainable parameters. ||||| | denotes concatenation along the feature dimension. $\mathbf{Z}\in\mathbb{R}^{N\times d_{LLM}}$ represents preliminary spatial tokens. $d_{LLM}$ is the token dimension of the LLM. $\bm{Z}_{mask}\in\mathbb{R}^{N\times d_{LLM}}$ indicates the mask tokens which contain the missing pattern of the data. $\bm{Z}_{S}\in\mathbb{R}^{N\times d_{LLM}}$ represents our target spatial tokens.
Temporal Tokenizer. The temporal tokenizer is designed to activate the LLM to understand spatial-temporal data by converting the spatial-temporal data into sequential tokens along the temporal dimension. We aim to encode the spatial-temporal data at each time slice into a token. Specifically, We first concatenate the spatial-temporal data, the time embedding, topology-aware node embedding, and the binary mask tensor, then aggregate various types of information through MLP to obtain the hidden representation $\mathbf{F}\in\mathbb{R}^{T\times N\times d_{T}}$. Subsequently, we utilize attention mechanism to aggregate the information of nodes in the hidden space, resulting in $\mathbf{H}_{T}\in\mathbb{R}^{T\times 1\times d_{T}}$. In this process, we need to aggregate information across the spatial dimension while ensuring that the model can adapt to different graph structures. Therefore, using methods such as self-attention is not suitable. To address this, we learn reference points $\mathbf{Q}_{l}\in\mathbb{R}^{T\times 1\times d_{T}}$ for each time step to aggregate information across the spatial dimension by interacting with nodes. The process is defined as follows:
@@
[…]
$\mathbf{H}_{T}=\text{Attention}(\mathbf{Q}_{l},\mathbf{F},\mathbf{F}),$
$\bm{Z}_{T}=\text{LayerNorm}(\mathbf{W}_{T}\mathbf{H}_{T}+\mathbf{b}_{T}),$
where $\mathbf{W}_{*}$ and $\mathbf{b}_{*}$ denote the trainable parameters. $d_{k}$ is the feature dimension, which is used for normalizing the variance in the attention mechanism. The MLP consists of two layers of linear with a ReLU activation function in between. Within the attention operation, we compute the similarity between the reference node and the hidden representation of each node, and subsequently aggregate the hidden representations of the nodes based on these similarities to obtain $\mathbf{H}_{T}$. Finally, a linear layer followed by LayerNorm is applied to obtain the temporal tokens $\bm{Z}_{T}\in\mathbb{R}^{T\times d_{LLM}}$.
Hypergraph Learning We can capture the pair-wise correlation using spatial tokens aggregated by nodes, but it is difficult to capture important non-pairwise and higher-order spatial-temporal dependencies. To capture higher-order spatial-temporal correlations, we project node-level data into a region-level. For spatial-temporal data, the region is an appropriate choice, as the influencing factors, such as human activities and weather geological changes, exhibit distinct regional characteristics. Therefore, we design an autoencoder module to learn the mapping relationships between nodes and regions. The encoder, hypergraph encode, projects spatial tokens into higher-order region-level tokens. The decoder, hypergraph decode, is responsible for decoding the regional latent vectors back into node-level latent vectors. A simple implementation is to construct the encoder and decoder based on prior knowledge, but this approach overlooks the dynamics of the mapping relationships and cannot be easily transferred between different graph structures. Thus, we opted for attention mechanism-based encoders and decoders to dynamically learn these mapping relationships from the data. The process of hypergraph encode constructing region-level tokens $\bm{Z}_{H}\in\mathbb{R}^{M\times d_{LLM}}$ ($M$ represents the number of regions) from spatial tokens is as follows:
$\mathbf{H}_{N}=\mathbf{W}_{N}\bm{Z}_{S}+\mathbf{b}_{N},$ (7)
$\bm{Z}_{H}=\text{LayerNorm}(\mathbf{W}_{H}\text{Attention}(\mathbf{H}_{l},\mathbf{H}_{N},\mathbf{H}_{N})+\mathbf{b}_{H}),$
where $\mathbf{H}_{N}\in\mathbb{R}^{N\times d_{H}}$ represents the lower-dimensional representation of $\bm{Z}_{S}$ after passing through a linear layer. $d_{H}$ is the hidden dimension. $\mathbf{H}_{l}\in\mathbb{R}^{M\times d_{H}}$ is a set of learnable parameters that extract regional-level representations.
After $\bm{Z}_{H}$ is processed by the LLM to obtain its hidden representation $\mathcal{Z}_{H}\in\mathbb{R}^{M\times d_{LLM}}$, it is necessary to map the regional-level hidden representation back to the node level by hypergraph decode to perform subsequent tasks:
$\mathbf{H}_{R}=\mathbf{W}_{hidden}\mathcal{Z}_{H}+\mathbf{b}_{hidden},$ (8)
$\mathcal{Z}_{N}=\text{LayerNorm}(\mathbf{W}_{R}\text{Attention}(\mathbf{H}_{N},\mathbf{H}_{l},\mathbf{H}_{R})+\mathbf{b}_{R}).$
@@
In this recovery process, $\mathbf{H}_{R}\in\mathbb{R}^{M\times d_{H}}$ is the lower-dimensional representation of $\mathcal{Z}_{H}$. $\mathcal{Z}_{N}\in\mathbb{R}^{N\times d_{LLM}}$ refers to the recovered node-level hidden representation from the LLM.
To ensure that the regional-level tokens constructed by the Hypergraph module can preserve the connectivity between nodes, a constraint loss function is designed. Furthermore, to prevent nodes from being ignored during the attention mechanism if their weights are too low, we have also added regularization term for the constraint loss. Assuming $S\in\mathbb{R}^{M\times N}$ represents the attention weights between each regional-level token and node token, the constraint loss $\mathcal{L}_{C}$ can be expressed as:
$\mathcal{L}_{C}=-\sum_{m=1}^{M}\sum_{i,j,(i\neq j)}S_{m,i}S_{m,j}A_{i,j}+\underbrace{\log\pi(\text{Softmax}(\sum_{m=1}^{M}S_{m,:})|{\bm{\alpha}})}_{regularization\ term}.$ (9)
@@
Here, $\pi$ is the Dirichlet distribution. ${\bm{\alpha}}$ is the parameter of the Dirichlet distribution. $S_{m,:}$ denotes the vector of attention weights for the $m$-th regional token across all node tokens, and its sum is fixed, which can encourage the regional tokens to aggregate information from node tokens that are connected in the graph. A thorough explanation is available in the Appendix E.
Virtual Node. The spatial-temporal data we encounter often does not exist within an isolated and closed system, where external changes can have a substantial impact. For example, in traffic flow forecasting tasks, the entry and exit of external vehicles can significantly affect it. Thus, it is necessary to model the impact of external systems. For spatial-temporal data, the impact is primarily manifested in the interplay between time and space. So we utilize an attention mechanism between $\mathbf{E}_{T}$ and $\mathbf{E}_{N}$ to obtain $T$ tokens that represent the impact of external systems. Additionally, a learnable token is introduced to model some inherent characteristics of the external system that are independent of time. By combining these two components, the final virtual node $\bm{Z}_{E}\in\mathbb{R}^{(T+1)\times d_{LLM}}$ is derived to represent the influence of external changes.
Versatility for Spatial-Temporal Forecasting and Imputation Tasks. The main differences between forecasting and imputation lie in missing values and context. Forecasting tasks typically have few missing values and can only forecast based on historical values. Imputation tasks, on the other hand, have more missing data and can impute based on both historical and future values. In the Equation 4, we have already enhanced the model’s perception of missing values through the mask tokens. For the differences in context, the generalization ability of LLM can also effectively address it. Therefore, STD-LLM possess the capability to unify these two tasks. Switching between forecasting and imputation tasks can be achieved by modifying the target labels in the training dataset. Additionally, forecasting can be viewed as imputation of future missing values, thus unifying the two tasks. The model’s processing flow can be summarized as follows:
$\mathcal{Z}=\text{LLM}([\bm{Z}_{E}||\bm{Z}_{T}||\bm{Z}_{H}]),$ (10)
$\mathcal{Z}_{H}=\mathcal{Z}[-M:,:],\quad\mathcal{Z}_{N}\leftarrow\mathcal{Z}_{H},$
$\mathbf{Y}={\rm MLP}([\mathbf{W}_{Y}\mathcal{Z}_{N}+\mathbf{b}_{Y}||\mathbf{X}]),$
where, $\mathbf{Y}\in\mathbb{R}^{N\times(P\times C)}$ is the forecasting or imputation result, $P$ is the number of target time steps. $\mathbf{X}$ is reshaped to $\mathbf{X}\in\mathbb{R}^{N\times(T\times C)}$ for processing convenience. $\mathbf{W}_{*}$ and $\mathbf{b}_{*}$ denote the trainable parameters of the linear layers. The "LLM" in equation refers to the main body of the pre-trained LLM. The last line of the formula denote to the decoding step of the Decoder.

### v2

@@ 3.2 Model Structure
Periodic-aware time embedding: Spatial-temporal data exhibit periodicity, which is essential for the forecasting and imputation task. To avoid overfitting and underfitting, we select two kinds of periodicity that are moderately coarse including time-of-day and day-of-week to create embedding dictionaries, $\mathbf{D}_{t}\in\mathbb{R}^{288\times d_{t}}$ and $\mathbf{D}_{w}\in\mathbb{R}^{7\times d_{t}}$, where $d_{t}$ is the dimension of the time embedding. By looking-up and concatenation along with broadcast operations, we finally obtain the time embedding $\mathbf{E}_{T}\in\mathbb{R}^{T\times N\times 2d_{t}}$.
Figure 1: Model architecture. Module A provides spatial-temporal embeddings for node and temporal information. The spatial-temporal tokenizers B construct temporal and spatial tokens from different perspectives. Module C builds region-level spatial tokens based on the node-level spatial tokens which is the output of B-2. Module D projects the hidden representations to target output. For the PLM, We utilize partial freezing and LoRA to fine-tune the multi-head attention, position embedding and layernorm layers.
3.2.2 Spatial-Temporal Tokenizer.
Utilizing a PLM to process spatial-temporal data, the most crucial step is to transform the spatial-temporal data into tokens that the PLM can handle, as all subsequent processing is based on the PLM’s capture of the relationships between these tokens. To this end, we meticulously introduce a spatial-temporal tokenizer to generate tokens from both the spatial and temporal dimensions. Specifically, each node generates a spatial token meanwhile two kinds of temporal tokens are designed to represent the current state and evolving pattern of the spatial-temporal data. Through employing the spatial-temporal tokenizer, the original spatial-temporal graph data is converted into token sequences, enabling the subsequently applied PLM not only to capture the temporal and spatial correlations whithin each kind of tokens but also to capture intertwined spatial-temporal correlations between the spatial and temporal tokens.
Spatial Tokenizer: The spatial tokenizer’s objective is to aggregate the temporal information of each node separately, generating a token that represents the state of each node, thereby enabling the PLM to model the complex spatial correlations between different nodes. For the state of each node, we believe it can be decomposed into a relatively static intrinsic state and a frequently changing dynamic state. The intrinsic state, determined by the node’s spatial structure and periodicity, captures the macroscopic characteristics of the node. The dynamic state, derived from historical data, captures the microscopic variations of the node.
Based on the above analysis, we model the intrinsic states $\bm{Z}_{\text{intrinsic}}\in\mathbb{R}^{N\times d_{PLM}}$ using node embeddings and time embeddings, while using historical data to model the dynamic state $\bm{Z}_{\text{dynamic}}\in\mathbb{R}^{N\times d_{PLM}}$. To unify the spatial-temporal prediction and imputation tasks, enabling the PLM to detect the patterns of missing is essential. Thus, we further construct mask tokens $\bm{Z}_{\text{mask}}\in\mathbb{R}^{N\times d_{PLM}}$ based on the mask matrix $\bm{\mathcal{M}}$. By combining the three elements, we obtain the target node-level spatial tokens $\bm{Z}_{S}\in\mathbb{R}^{N\times d_{PLM}}$. Formally,
$\bm{Z}_{\text{intrinsic}}=\text{MLP}([\mathbf{E}_{T}||\mathbf{E}_{N}]),$ (3)
$\bm{Z}_{\text{dynamic}}=\text{MLP}(\mathcal{X}),$
$\bm{Z}_{\text{mask}}=\text{MLP}(\bm{\mathcal{M}}),$
$\bm{Z}_{S}=\text{LayerNorm}(\bm{Z}_{\text{dynamic}}+\bm{Z}_{\text{intrinsic}}+\bm{Z}_{\text{mask}}).$
In this process, we integrate the temporal dimension into the feature dimension. The MLP consists of two layers of linear with a ReLU activation function in between. ||||| | denotes concatenation along the feature dimension. $d_{PLM}$ is the token dimension of the PLM.
Temporal Tokenizer: Relying solely on spatial tokens can effectively model the spatial-temporal relationships between nodes, but the model may lack an understanding of the overall state and changing trends of the system. Therefore, we design temporal tokenizer that aggregates information from all the nodes for each time step to encapsulate the overall state and changing trends. Considering that a single time step cannot capture the sequence’s locality and trends, which are essential pieces of information, we draw on the concept from PatchTST  [42], merging all time steps into a single patch. Consequently, only one overall state token $\bm{Z}_{\text{state}}\in\mathbb{R}^{1\times d_{PLM}}$ and one overall trend token $\bm{Z}_{\text{trend}}\in\mathbb{R}^{1\times d_{PLM}}$ are produced, rather than generating separate tokens for each time step.
Specifically, we first take the mean of all node to obtain the overall average state $\bm{\mathcal{\overline{X}}}\in\mathbb{R}^{1\times(T\times C)}$. Then, we use the first difference of $\bm{\mathcal{\overline{X}}}$, denoted as $\bm{\mathcal{\overline{X}}}_{\text{trend}}\in\mathbb{R}^{1\times((T-1)\times C)}$, to represent the overall trend. Finally, $\bm{\mathcal{\overline{X}}}$ and $\bm{\mathcal{\overline{X}}}_{\text{trend}}$ are respectively combined with the time embeddings. A MLP is further utilized to perform feature transformation, resulting in $\bm{Z}_{\text{state}}$and $\bm{Z}_{\text{trend}}$. Then we concatenate these two tokens and apply Layer Normalization to normalize them, thereby obtaining the temporal tokens $\bm{Z}_{T}\in\mathbb{R}^{2\times d_{PLM}}$. The above process is defined as follows:
$\bm{Z}_{\text{state}}=\text{MLP}(\bm{\mathcal{\overline{X}}}||\mathbf{E}_{T}[T-1:,0,:]),$ (4)
$\bm{Z}_{\text{trend}}=\text{MLP}(\bm{\mathcal{\overline{X}}}_{\text{trend}}||\mathbf{E}_{T}[T-1:,0,:]),$
$\bm{Z}_{T}=\text{LayerNorm}([\bm{Z}_{\text{state}}||\bm{Z}_{\text{trend}}]),$
where the $\mathbf{E}_{T}[T-1:,0,:]\in\mathbb{R}^{1\times d_{t}}$ represents the time embedding of the last time step in the input.
3.2.3 Sandglass Attention.
Taking into account the following two factors, we have developed a sandglass attention module. 1) The number of spatial tokens $N$ is usually large in real-world applications which significantly hampers model efficiency especially the training and inference speed. 2) The spatial tokens primarily encapsulate information at the node level. However, these node-level spatial tokens alone struggle to capture important non-pairwise and higher-order spatial-temporal correlations. Specifically, the proposed sandglass attention (SGA) module first aggregates node-level spatial tokens into fewer region-level spatial tokens thoughr a precoder, enabling the model to capture non-pairwise and higher-order spatial-temporal correlations while enhancing computational efficiency. Then a decoder is employed to restore the original length of spatial tokens. Next, we detail the SGA module.
In the SGA precoder, we aggregate the node-level spatial tokens $\bm{Z}_{S}$ into region-level spatial tokens $\bm{Z}_{H}\in\mathbb{R}^{M\times d_{LLM}}$ through a learnable query matrix $\mathbf{H}_{l}\in\mathbb{R}^{M\times d_{PLM}}$:
$\bm{Z}_{H}=\text{LayerNorm}((\text{Attention}(\mathbf{H}_{l},\bm{Z}_{S},\bm{Z}_{S})).$ (5)
where $M$ represents the number of region-level spatial tokens, and $M<N$. Besides, the scaled dot-product-attention is employed in our paper. Formally,
@@
We then feed $\bm{Z}_{H}$ into PLM to further capture the spatial-temporal correlations between the region-level spatial tokens and temporal tokens. Finally, we get the corresponding $\bm{Z}_{H}^{{}^{\prime}}\in\mathbb{R}^{M\times d_{PLM}}$. To fit with the output size, it is necessary to map the region-level hidden representation back to the node-level by a SGA decoder:
$\bm{Z}_{N}^{{}^{\prime}}=\text{LayerNorm}((\text{Attention}(\bm{Z}_{S},\mathbf{H}_{l},\bm{Z}^{{}^{\prime}}_{H})),$ (7)
@@
where $\bm{Z}_{N}^{{}^{\prime}}\in\mathbb{R}^{N\times d_{PLM}}$ refers to the recovered node-level hidden representation from the PLM.
3.2.4 Unified Output Projection.
Given that both forecasting and imputation tasks fundamentally involve understanding the complex spatial-temporal dependencies of data, and that the PLM has completed a substantial amount of imputation tasks during pre-train phase, we expect our model, which is designed based on PLM, can handle both of the forecasting and imputation tasks. Thus, we need to enable our model to recognize the presence of missing values. In the Eq. 3, we have already enhanced the model’s perception of missing values through the mask tokens. Therefore, in the output layer, we unified these two task based on the following projection:
$[\bm{Z}_{T}^{{}^{\prime}},\bm{Z}_{H}^{{}^{\prime}}]=\text{PLM}([\bm{Z}_{T}||\bm{Z}_{H}]),\quad\bm{Z}_{N}^{{}^{\prime}}\leftarrow\bm{Z}_{H}^{{}^{\prime}},$ (8)
$\mathbf{Y}={\rm MLP}(\bm{Z}_{N}^{{}^{\prime}}+\bm{Z}_{T}^{{}^{\prime}}[1:,:]+\bm{Z}_{S}),$
where $\mathbf{Y}\in\mathbb{R}^{N\times(T\times C)}$ is the forecasting or imputation result. $\bm{Z}_{T}^{{}^{\prime}}\in\mathbb{R}^{2\times d_{PLM}}$ is the hidden representations of temporal tokens. After constructing the tokens and inputting them into the PLM, we obtain the hidden representations. Then, we sum the hidden representation of ecah node with the hidden representation of the overall trend temporal token, and make a residual connection with the node-level spatial tokens to obtain the hidden representation of the target state. Finally, a MLP is used to map the target state hidden representations to our target outputs.
3.2.5 Model Training.
Given that the PLM has only been exposed to textual data with a significantly different distribution during the pre-training phase and lacks knowledge related to spatial-temporal data, it is necessary to fine-tune the PLM to acquire knowledge of spatial-temporal data and adapt to the distribution patterns of such data. Furthermore, considering that the SGA module needs to learn a query matrix from scratch, this process may lead to overfitting and confusion in spatial correlation if no constraints such as the adjacency matrix are introduced, resulting in a decrease in model performance. Therefore, we designed a constrained loss function to guide and constrain the training of the SGA module. The specific design is detailed in the following text.
Fine-tuning PLM: Based on some existing studies [7, 37], we realized that the knowledge of the PLM is primarily stored in the attention modules, while layer normalization and position embeddings are sensitive to the distribution of the data. Therefore, we choose to fine-tune the attention layers to embed spatial-temporal knowledge into the PLM, while also fine-tuning the position embeddings and layer normalization to adapt the PLM to the distribution of spatial-temporal data. To reduce the number of parameters needing updates, we apply Low-Rank Adaptation (LoRA) [43] to the attention layers. Meanwhile, the position embeddings and layer normalization, with fewer parameters, are fully updated.
Loss Function: We use the L1 loss between the model’s output $Y$ and the ground truth to train our model. Utilizing only this loss function can achieve satisfactory results on the training set. However, to enable the model to make more effective use of the graph structure and to mitigate the phenomenon of overfitting, further design is required. Specifically, to ensure that in the SGA module, the learned region-level spatial nodes can reflect the original graph structure $G$, we design a structure-aware loss function $\mathcal{L}_{G}$ when aggregating node-level spatial tokens. Furthermore, to prevent overfitting that might cause some node information to be ignored (attention weights are close to zero), we introduce a regularization term $\mathcal{L}_{R}$. The sum of these two components forms constraint loss function $\mathcal{L}_{C}=\mathcal{L}_{G}+\mathcal{L}_{R}$. Assuming $S\in\mathbb{R}^{M\times N}$ represents the attention weights between each region-level spatial token and node-level spatial token, $\mathcal{L}_{G}$ and $\mathcal{L}_{R}$ can be expressed as:
$\mathcal{L}_{G}=-\sum_{m=1}^{M}\sum_{i,j(i\neq j)}{S}_{m,i}{S}_{m,j}A_{i,j},$ (9)
$\mathcal{L}_{R}=-\log\pi(\text{Softmax}(\sum_{m=1}^{M}S_{m,:})|\bm{\alpha}),$
@@
where $\pi$ is the Dirichlet distribution. ${\bm{\alpha}}\in\mathbb{R}^{N}$ is the parameter of the Dirichlet distribution. $S_{m,:}\in\mathbb{R}^{N}$ denotes the vector of attention weights for the $m$-th region token across all node tokens, which satisfies the properties: $\sum_{i=1}^{N}S_{m,i}=1$.
It is clear that the more edges there are, the smaller the minimum value of $L_{G}$ can be. Consider a complete graph of size P, where:
$\mathcal{L}_{G}$ $=-\sum_{m=1}^{M}\sum_{i,j(i\neq j)}{S}_{m,i}{S}_{m,j}\cdot 1,$ (10)
$=-\sum_{m=1}^{M}\sum_{i}{S}_{m,i}(1-{S}_{m,i})=-M+\|{S}\odot{S}\|_{1}.$
The $\odot$ denotes the Hadamard Product. It is evident that minimizing the objective function $\mathcal{L}_{G}$ is achieved by evenly distributing the values of $S$, such that $S_{m,i}=\frac{1}{P}$. Under this condition, $\mathcal{L}_{G}$ simplifies to $-\frac{M(P-1)}{P}$. Shifting the discussion to a regular graph of order $N$, a lower $\mathcal{L}_{G}$ facilitates the SGA module’s ability to concentrate on larger complete subgraphs or those with robust connectivity, effectively integrating the graph’s structural properties into the model.
The essence of $\mathcal{L}_{R}$ is simple: the expected value of $\pi$ is $\bm{\alpha}/{\sum_{i}\bm{\alpha}_{i}}$, indicating that a larger $-\mathcal{L}_{R}$ drives $\text{Softmax}(\sum_{m=1}^{M}S_{m,:})$ closer to this expected value. By tuning $\bm{\alpha}$, we indirectly control the attention weights distribution. Setting $\bm{\alpha}=\{1.05\}^{N}$ (a small value greater than 1) ensures equal aggregation of nodes. Yet, to prioritize significant nodes, we refine $\bm{\alpha}$ by adding the node degrees: $\bm{\alpha}=\{1.05\}^{N}+\text{Softmax}(\text{diag}(D))$.